# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v4

@@ 4.1 Maximum likelihood estimation and ergodic stationarity
The parameters of the structural STVAR models discussed in Section 3 can be estimated by the method of maximum likelihood (ML), which amounts to maximizing the log-likelihood function over the parameter space. To obtain a well-behaving estimator with desirable asymptotic properties such as consistency, the parameter space is often restricted to the region where the model is ergodic stationary. Based on the results of Saikkonen (2008) (see also Kheifets and Saikkonen, 2020), we show in Appendix A (Theorem A.1) that when the transition weights $\alpha_{m,t}$ are $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one. But since this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not very practical to use it to restrict the parameter space in estimation. Therefore, we restrict the parameter space with the more easily verifiable condition that the usual stability condition is satisfied for each of the regimes (Condition A.1 in Appendix A), as it is necessary for the sufficient condition. After estimation, the sufficient condition can be checked for the solutions of interest.
Maximizing the log-likelihood function (presented in Section 4.2) can be challenging in practice due to its high multimodality, induced by the nonlinear dynamics. Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly and are not often even able to converge to a local maximum. To overcome this issue, we propose to allow for unstable estimates and to maximize the penalized log-likelihood function, obtained by adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close to the boundary of the stability region. In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance, while sufficient penalization eventually steers the optimization algorithm back to the stability region.
Penalized likelihood -based estimation has been previously applied to time series models by Nielsen and Rahbek (2024). They impose nonnegativity constraints in the autoregressive conditional heteroscedasticity model, allowing the estimation algorithm to explore the boundary of the parameter space. Nielsen and Rahbek (2024) use penalization also to determine which parameters should be set to zero.
@@ 4.2 The penalized log-likelihood function
Before introducing the penalized log-likelihood function, the log-likelihood function is presented. We assume that each shock $e_{it}$, $i=1,...,d$, follow the skewed $t$-distribution introduced by Hansen (1994) with zero mean, unit variance, $\nu_{i}>2$ degrees of freedom, and skewness controlled by the parameter $\lambda_{i}\in(-1,1)$. The density function of the skewed $t$-distribution, $st(\cdot;\nu_{i},\lambda_{i})$, is given as (Hansen, 1994, Equations (10)-(13)):
@@
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is defined in Section 2 and $B_{y,t}$ in Equation (3.3). The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{M},$ $\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degree-of-freedom and skewness parameters. [^2] With threshold (2.3) or logistic (2.4) weights, $\alpha=(r_{1},...,r_{M-1})$ or $\alpha=(c,\gamma)$, respectively, whereas with exogenous weights, $\alpha$ is omitted from the parameter vector.
The penalized likelihood function is then defined as
@@ 4.3 Three-step estimation procedure
Lanne et al. (2017) propose a three-step estimation procedure for their linear SVAR model identified by non-Gaussianity. In the first step, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent steps to maximize the log-likelihood function. The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a high number of modes to the (penalized) log-likelihood function. To address these challenges, we propose a modified version of the three-step estimation procedure in which multimodality is explicitly taken into account.
@@
    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1},...,\phi_{M},\varphi_{1},...,\varphi_{M},\alpha)$ by the method of (penalized) nonlinear least squares (NLS) (see Appendix C for details). Denote this estimate of $\boldsymbol{\theta}_{1}$ as $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
@@
Step 1 obtains initial estimates for the AR and weight function parameters, and it is comparable to Step 1 of Lanne et al. (2017) but with the nonlinearity of the model explicitly accounted for. Notably, the NLS estimation often produces estimates that do not satisfy the usual stability condition for each of the regimes, and thus allowing for instability (but penalizing it) facilitates utilization of the NLS estimates in our three-step procedure. Step 2 is comparable to Step 2 of Lanne et al. (2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able to escape from local maxima. Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps, often converging to a nearby local maximum of the (penalized) log-likelihood function.
@@
Due to the high multimodality of the (penalized) log-likelihood function, Steps 2 and 3 should be ran a large number times to improve the reliability of the results. Since multiple estimation rounds are run, the estimation procedure produces a set of estimates, presumably corresponding to different local maximums of the (penalized) log-likelihood function. If there are multiple solutions with (penalized) log-likelihoods almost equal to the greatest found (penalized) log-likelihood, the blended identification procedure discussed in Section 3.5 can be employed to obtain the best local solution that facilitates labelling the shocks of interest. The proposed three-step estimation procedure is implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
@@
To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-step estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix D. Since estimation is computationally demanding, we consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$. To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both TVAR and LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive.
According to the results (see Tables D.2 and D.3 in Appendix D for details), estimation accuracy is slightly better when the AR matrices are inside the stability region, but even when they are in the penalization region, the estimators of the AR matrices exhibit only minor small-sample bias. Additionally, estimation accuracy appears to be slightly higher for the TVAR models compared to the LSTVAR models. While for most of the parameters there is only very slight bias, we find a notable upward small-sample bias for the scale parameter, for the degrees-of-freedom parameter whose true value is not small, and for the first entry in each impact matrix. The upward biases in the scale and not-small degrees-of-freedom parameters can both be explained by weak discrimination between large and very large estimates, as large values already imply nearly discrete regime-switches and near-Gaussian distributions, respectively. The upward bias in the first entry of each impact matrix, in turn, could be explained by its sensitivity to extreme values induced by the related particularly heavy tailed skewed $t$-distribution. Nonetheless, the bias diminishes for all of the parameters as the sample size increases. Since also the standard deviations of the estimates diminish as the sample size increases, our results are in line with consistency of the PML estimator. Overall, the performance of the proposed PML estimator and the three-step estimation procedure seems reasonable.
@@ 5 Empirical application
Our empirical application studies the macroeconomic effects of climate policy uncertainty (CPU). In the related literature, Fried et al. (2022) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy transition risk shifts investments toward cleaner capital while reducing total investments and output. Khalil and Strobel (2023) reach similar conclusions in their dynamic stochastic general equilibrium (DSGE) model, attributing the effects to financial institutions’ aversion to uncertainty. Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, reducing output and increasing inflation, which is supported by results from their recursively identified Bayesian VAR.
We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty (EPU). Such variation can be accommodated in our structural STVAR model by specifying a logistic transition weight function with the (lagged) level of EPU as the switching variable. This approach allows us to capture potential nonlinearities and regime-dependent effects, providing more detailed results on the macroeconomic effects of the CPU shock than linear SVAR analysis.
@@ 5.1 Data
We consider a monthly U.S. dataset consisting of six variables and covering the time period from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure CPU with the climate policy uncertainty index (CPUI) of Gavriilidis (2021), which is constructed based on the amount of newspaper coverage on climate policy uncertainty topics. The availability of the CPUI data determines the beginning of our sample period. In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) (Baker et al., 2016), which is based on newspaper coverage on topics related to economic policy uncertainty as well as on components quantifying the present value of future scheduled tax code expirations and disagreement among professional forecasters over future government purchases and consumer prices.
As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking first differences. For measuring the price level, we use the logs of consumer price index (CPI) and producer price index (PPI), likewise detrended by taking first differences. Finally, we measure the monetary policy stance with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE). The series of the included variables are presented in Figure 1.333The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, PPI, and the Federal funds rate are retrieved from the Federal reserve bank of St. Louis database and the Wu and Xia (2016) shadow rate from the Federal reserve bank of Atlanta’s website.
Figure 1: Monthly U.S. time series covering the period from 1987:4 to 2024:12. From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The bottom panel shows the transition weights of the High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model. The shaded areas indicate the U.S. recessions defined by the NBER.
@@ 5.2 The structural LSTVAR model
In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks must be specified. We employ the logistic transition weights with two regimes (Anderson and Vahid, 1998) given in Equation (2.4). They facilitate smooth transitions between the regimes and enable the analysis of how the effects of the CPU shock vary depending on the level of EPU, whose first lag is specified as the switching variable. The logistic weight function also has the convenient property that is nests the threshold weight function (Tsay, 1998) as a special case when the scale parameter tends to infinity, effectively incorporating the selection between these two weight functions within estimation.
We estimate the structural LSTVAR model by PML with the three-step procedure proposed in Section 4, assuming that the structural shocks follow the independent skewed $t$-distributions. The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the validity of the identifying Assumption 1 can be assessed based on the estimates of these parameters. To avoid overfitting, particularly in the regime accommodating fewer observations, the autoregressive order $p=2$ is selected based on Hannan-Quinn information criterion.
To deal with the problem of weak identification with respect to the ordering and signs of the columns of $B_{2}$, discussed in Section 3.3, 3.4, and 3.5, we implement the following, simplistic blended identification strategy that largely avoids making restrictive assumptions about the effects of the CPU shock. We normalize the first row of $B_{1}$ positive with CPUI ordered first and assume that in both regimes, one of the shocks, deemed as the CPU shock, has greater impact effect on the CPUI than any other shock, and this impact effect has the same sign in both regimes. Then, to support our identification, we check that the recovered CPU shock aligns with several major historical events affecting CPU. In particular, we find that President Trump’s election is preceded by large CPU shocks in October 2016 and 2024 ($17$ and $26$, respectively). The largest negative recovered CPU shock ($-10$), in turn, dates to August 2022, when the Inflation Reduction Act was passed, providing support for the clean energy transition. Among all the shocks recovered from the model, only the one deemed as the CPU shock has signs consistent with the above historical events affecting CPU.
@@
The location parameter estimate is $\hat{c}=14.48$, implying roughly that Regime 1 dominates when the level of EPU is not high, while Regime 2 starts prevailing in times of high EPU. Therefore, we label Regime 1 as Low EPU Regime and Regime 2 as High EPU Regime. The scale parameter estimate is $\hat{\gamma}=16.65$, indicating that transitions between the regimes are somewhat fast (rather than very smooth), which can be observed also from the bottom panel of Figure 1, where the fitted transition weights of High EPU Regime are presented.
@@
The estimates of the degrees-of-freedom and skewness parameters are (2.33,2.10,2.64,3.23,6.54,(2.33,2.10,2.64,3.23,6.54,( 2.33 , 2.10 , 2.64 , 3.23 , 6.54 , 2.68)2.68)2.68 ) and $(0.41,0.17,0.09,0.43,0.25,-0.12)$, respectively. Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied. Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes. [^4] We compute the upper bound using the JSR toolbox (Jungers, 2023) in MATLAB. The obtained upper bound is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1. Based on the residual diagnostics presented in Appendix E, the adequacy of our model to capture the autocorrelation structure, conditional heteroskedasticity, and distribution of the data seems, overall, reasonable.
5.3 Generalized impulse response functions

### v5

@@ 4.1 Maximum likelihood estimation and ergodic stationarity
The parameters of the structural STVAR model discussed in Section 3 can be estimated by the method of maximum likelihood (ML). To obtain a well-behaving estimator with desirable asymptotic properties such as consistency, the parameter space is often restricted to the region where the model is ergodic stationary. Building on the results of Saikkonen (2008) (see also Kheifets and Saikkonen, 2020), it can be shown that when the transition weights $\alpha_{m,t}$ are logistic (2.3) or of the threshold form, a sufficient condition for ergodic stationarity is that the joint spectral radius of the companion form AR matrices of the regimes is strictly less than one (see Theorem B.1 in Appendix B).555 For logistic models (with $M=2$ assumed), we additionally require that the matrix $B_{1}^{-1}B_{2}$ has no negative real eigenvalues to rule out singular convex combinations of $B_{1}$ and $B_{2}$ (see Appendices B and C.5 for details). See Appendix A for the definition of the threshold weights. Because this condition is computationally demanding to verify (e.g., Chang and Blondel, 2013), it is not particularly useful for restricting the parameter space during estimation. Therefore, we instead impose the usual stability condition for each regime (Condition B.1 in Appendix B), as it is a necessary condition for the sufficient one. The sufficient condition can then be checked after estimation for the solutions of interest.
Maximizing the log-likelihood function can be challenging in practice due to its high multimodality, induced by the nonlinear dynamics. [^6] The log-likelihood function is presented in Section 4.2 for shocks that follow skewed $t$-distributions. Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly. To overcome this issue, we propose to allow for unstable estimates and to maximize the penalized log-likelihood function, obtained by adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close to the boundary of the stability region. In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance, while sufficient penalization eventually steers the algorithm back to the stability region. [^7] Penalized likelihood-based estimation has been previously applied to time series models by Nielsen and Rahbek (2024). They impose nonnegativity constraints in the autoregressive conditional heteroskedasticity model, allowing the estimation algorithm to explore the boundary of the parameter space. Nielsen and Rahbek (2024) use penalization also to determine which parameters should be set to zero.
@@ 4.2 The penalized log-likelihood function
Before introducing the penalized log-likelihood function, the log-likelihood function is presented. We assume that each shock $e_{it}$, i=1,,di=1,...,d, follows the skewed $t$-distribution introduced by Hansen (1994) with zero mean, unit variance, $\nu_{i}>2$ degrees of freedom, and skewness controlled by the parameter $\lambda_{i}\in(-1,1)$. The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the plausibility of the identifying Assumption 1 can be (informally) assessed based on the estimates of these parameters.
The density function of the skewed $t$-distribution, $st(\cdot;\nu_{i},\lambda_{i})$, is given as (Hansen, 1994, Equations (10)-(13)):
@@
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is the conditional mean defined in Section 2 and $B_{y,t}$ is the impact matrix defined in Equation (3.3). The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{M},\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),$..., $\text{vec}(A_{m,p}))$, m=1,,Mm=1,...,M, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degrees-of-freedom and skewness parameters. With logistic weights (2.3), $\alpha=(c,\gamma)$, whereas with exogenous weights, $\alpha$ is omitted from the parameter vector.
The penalized log-likelihood function is then defined as
@@ 4.3 Three-step estimation procedure
Lanne et al. (2017) propose a three-step estimation procedure for their linear SVAR model identified by non-Gaussianity. In the first step, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent steps to maximize the log-likelihood function. The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a large number of modes to the (penalized) log-likelihood function. To address these challenges, we propose a modified version of the three-step estimation procedure in which multimodality is explicitly taken into account.
@@
    Estimate the autoregressive and weight function parameters $\boldsymbol{\theta}_{1}\equiv(\phi_{1},...,\phi_{M},\varphi_{1},...,\varphi_{M},\alpha)$ by (penalized) nonlinear least squares (NLS) (see Appendix D for details). Denote this estimate of $\boldsymbol{\theta}_{1}$ as $\hat{\boldsymbol{\theta}}_{1}^{NLS}$.
@@
Step 1 obtains initial estimates for the AR and weight function parameters, and it is comparable to Step 1 of Lanne et al. (2017) but with the nonlinearity of the model explicitly accounted for. Notably, the NLS estimation often produces estimates that do not satisfy the usual stability condition for each of the regimes, and thus allowing for instability (but penalizing it) facilitates utilization of the NLS estimates in our three-step procedure. Step 2 is comparable to Step 2 of Lanne et al. (2017), but the multimodality of the (penalized) log-likelihood function is taken into account by making use of a robust estimation algorithm that is able to escape from local maxima. Step 3 finalizes the estimation by initializing a standard gradient based optimization algorithm from the initial estimates obtained from the previous steps.
@@
Due to the high multimodality of the (penalized) log-likelihood function, we recommend running Steps 2 and 3 a large number of times to improve the reliability of the results. Since a large number of estimation rounds are run, the estimation procedure produces a set of estimates, some of which presumably corresponding to different local maximums of the (penalized) log-likelihood function. If there are multiple solutions with (penalized) log-likelihoods close to the greatest found (penalized) log-likelihood, the blended identification strategy discussed in Section 3.3 can be employed to address weak identification.
@@
To assess the performance of the penalized maximum likelihood (PML) estimator as well as our three-step estimation procedure, we conduct a small-scale Monte Carlo study, which is discussed in detail in Appendix E. Since estimation is computationally demanding, we consider the simple bivariate (structural) LSTVAR model with $p=1$ and $M=2$. To evaluate how the PML estimator performs when the AR matrices are in the penalization region, we consider two specifications of both of LSTVAR models: one where the AR matrices are well inside the stability region and one where they lie clearly in the region where the penalization term is strictly positive. According to the results (see Table E.2 in Appendix E for details), the estimation accuracy is slightly better when the AR matrices are inside the stability region. Both specifications exhibit some bias in small samples. However, since both the bias and the standard deviations of the estimates diminish as the sample size increases, our results align with (possible) consistency of the PML estimator.
@@ 5 Empirical application
Our empirical application studies the macroeconomic effects of climate policy uncertainty (CPU). In the related literature, Fried et al. (2022) develop a dynamic general equilibrium model incorporating beliefs about future climate policy, and find that an increased climate policy transition risk shifts investments toward cleaner capital while reducing total investments and output. Khalil and Strobel (2023) reach similar conclusions in their dynamic stochastic general equilibrium (DSGE) model, attributing the effects to financial institutions’ aversion to uncertainty. Huang and Punzi (2024) also use a DSGE model and find that CPU shocks cause firms to delay investments, which reduces output and raises inflation. These findings are supported by their recursively identified Bayesian SVAR model.
We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty (EPU), as the level of uncertainty may affect the behavior of economic agents. Such variation can be accommodated in our structural STVAR model by specifying a logistic transition weight function with the (lagged) level of EPU as the switching variable. This approach allows us to capture potential state-dependent effects, providing more detailed results than linear SVAR analysis.
@@ 5.1 Data
We consider a monthly U.S. dataset consisting of five variables and covering the time period from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure CPU with the climate policy uncertainty index (CPUI) of Gavriilidis (2021), which is constructed based on the amount of newspaper coverage on climate policy uncertainty topics. In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) of Baker et al. (2016), which is based on newspaper coverage on topics related to economic policy uncertainty as well as on components quantifying the present value of future scheduled tax code expirations and disagreement among professional forecasters over future government purchases and consumer prices.
As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking first differences. For measuring the price level, we use the log of consumer price index (CPI), likewise detrended by taking first differences. Finally, the monetary policy stance is measured with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE). The series of the included variables are presented in Figure 1.888The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, and the Federal funds rate are retrieved from the Federal Reserve Bank of St. Louis database and the Wu and Xia (2016) shadow rate from the Federal Reserve Bank of Atlanta’s website. The availability of the CPUI data determines the beginning of our sample period.
Figure 1: Monthly U.S. time series covering the period from 1987:4 to 2024:12. From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index and consumer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The bottom panel shows the transition weights of the High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model. The shaded areas indicate the U.S. recessions defined by the NBER.
@@ 5.2 The structural LSTVAR model
The structural STVAR model is specified from (3.2)-(3.3) by defining the transition weights and distributions of the shocks. We employ the logistic transition weights with two regimes given in Equation (2.3). They facilitate smooth transitions between the regimes and enable analyzing how the effects of the CPU shock vary depending on the level of EPU, whose first lag is specified as the switching variable. We estimate the structural LSTVAR model by PML with the three-step procedure proposed in Section 4, and the shocks are assumed to follow skewed $t$-distributions. The autoregressive order $p=2$ is selected based on the Hannan-Quinn information criterion. The relatively low autoregressive order avoids overfitting, particularly in the regime accommodating fewer observations.
To deal with the problem of weak identification discussed in Section 3.2 and 3.3, we implement a blended identification strategy that imposes the following overidentifying restrictions on the impact matrices $B_{1}$ and $B_{2}$. First, for each variable, one of the shocks has a greater impact effect on that variable in both regimes than the other shocks, this shock is unique to that variable, and the impact effect has the same sign in both regimes. Second, the shock that has the greatest impact effect on output moves inflation in the same direction in both regimes (consistent with a demand shock). Third, the shock that has the greatest impact effect on inflation moves output in the opposite direction in both regimes (consistent with a supply shock). Fourth, the shock that has the greatest impact effect on interest rate moves output in the opposite direction in both regimes (consistent with a monetary policy shock). The shock that has the greatest impact effect on CPUI is deemed as the CPU shock, and to support our identification, we check that the recovered CPU shock aligns with several major historical events affecting CPU. In particular, we find that President Trump’s election is accompanied by large positive CPU shocks in November 2016 and 2024. Moreover, there is a large negative CPU shock in January 2021, when President Biden signed executive orders to re-enter the Paris Agreement and launch a broad federal climate strategy.
@@
The location parameter estimate $\hat{c}=16.16$, roughly implying that Regime 1 dominates when the level of EPU is not high, while Regime 2 prevails in times of high EPU. Hence, we label Regime 1 as the Low EPU Regime and Regime 2 as the High EPU Regime. The scale parameter estimate $\hat{\gamma}=64.23$, indicating that transitions between the regimes are quite fast, which can be observed also from the bottom panel of Figure 1, where the evolution of the fitted transition weight of the High EPU Regime is depicted.
@@
The estimated vectors of the degrees-of-freedom and skewness parameters are $(2.77,2.71,3.00,3.15,2.03)$ and (0.38,0.41,0.16,0.15,0.11)(0.38,0.41,-0.16,0.15,0.11), respectively, in the order of the variables CPUI, EPUI, IPI, CPI, and RATE. Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is plausible. Then, we check whether our model satisfies the stationarity condition by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes (see Theorem B.1 in Appendix B). The obtained upper bound is strictly less than one ($0.98$), implying that our model is ergodic stationary. [^9] In addition, we checked that the estimated $B_{1}^{-1}B_{2}$ does not have real negative eigenvalues, so the additional condition mentioned in Footnote 5 for logistic weights is satisfied. Finally, based on graphical residual diagnostics, the overall adequacy of our model seems reasonable (for details, see Appendix F).
5.3 Impulse response analysis